OpenStudy (anonymous):

Are these problems inconsistent? 2x+3y=2 2x-3y=6

OpenStudy (anonymous):

I'm not too sure on the difference between consistent and inconsistent equations.

OpenStudy (anonymous):

You typed before: Inconsistent -- no solution Consistent -- a single solution Dependent -- infinite number of solution This system has 1 solution. You can find it by elimination or substitution.

OpenStudy (anonymous):

okay. i see.

OpenStudy (anonymous):

Consistent -- a single solution Just add them and then subtract them.

OpenStudy (anonymous):

If the slope is different, then the two lines will intersect and will have one solution. They then form a Consistent and independent system with one solution. If the slope is same, y-intercept is same, then the two lines coincide. They then form a Consistent and dependent system with infinite solutions. If the Slope is same but different y-intercepts, then the two lines are parallel. They then form an Inconsistent system with no solution. 2x+3y=2 2x-3y=6 Putting them in slope-intercept form 2x+3y=2 3y = - 2x + 2 2 2 y = - ---- x + ----- 3 3 2x-3y=6 -3y = -2x + 6 -2 6 y = ---- x + ----- -3 -3 2 y = ---- x - 2 3 since both have different slopes, hence they will intersect and have one solution. so they form a consistent pair

OpenStudy (anonymous):

thanks guys!